# Uncertainty Calibration of Multi-Label Bird Sound Classifiers

### v1

@@ 4.2 Temperature and Platt Scaling
Results for global TSare mixed. While ConvNeXtBS and Perch v2 show improvements in MCS on all datasets ranging from 3.2% to 53%, AudioProtoPNet and BirdMAE exhibit more varied performance, with some datasets showing degradation in MCS by over 938%. On some datasets the improvements are substantial, with up to 98%. For AudioProtoPNet, TSseems to degrade calibration substantially on UHH and HSN and very slightly on PER. BirdMAE improves except for UHH. PSfollows the same trend as TS, with slightly better performance in most cases.
@@
The difference of MCS on the remaining test set (- first 10 min) compared to the full test set is marginal for all models. Comparing global and per-class scaling parameters is therefore valid. Per-class TSshows improvements for all models and datasets except for AudioProtoPNet on UHH POW, and HSN, BirdMAE on POW and UHH and Perch v2 on POW and SNE. Here, also substantial degradations of up to -288% are observed. Improvements range up to 97%. PSonly degrades on POW and PER for AudioProtoPNet and POW for Perch v2. Otherwise, improvements range up to 100% outperforming TSin most cases.
Both post hoc calibration methods demonstrate that substantial improvements to the deployment specific calibration can be achieved with minor computational requirements. Exemplary, AudioProroPNet appearing better calibrated than Perch v2 after per-class PS. However, no model or post hoc method consistently outperforms the others across all datasets, which reiterates that calibration and improvements using post hoc methods are domain-specific.
@@ 4.3 Implications for deployment
    When communicating uncertainty to end users (e.g., survey practitioners), reporting OCS/UCS alongside ECE or MCS helps reveal whether observed errors stem from systematic over or underconfidence and whether that profile differs between common vs. rare species.
@@ 5 CONCLUSIONS
We presented a first comprehensive benchmark of uncertainty calibration for four multi-label deep learning bird sound classifiers across eight expert-annotated BirdSet evaluation datasets. Beyond reporting class-wise mean average precision (cmAP), we evaluated calibration holistically using reliability diagrams and threshold-free metrics: expected calibration error (ECE), miscalibration score (MCS), and the decomposition into over and underconfidence (OCS/UCS). We analysed calibration at three granularities—global, per-dataset, and per-class—and examined behaviour for common versus rare species. Finally, we assessed lightweight post hoc methods (temperature and Platt scaling) fitting global calibration parameters on the POW validation set and per-class parameters on the first 10 min of each test datasets.
Three findings stand out. First, calibration varies substantially across datasets and classes; aggregating across domains can mask systematic miscalibration. In our setting, Perch v2 and ConvNeXtBS are consistently underconfident (near-zero OCS, UCS-dominated MCS), whereas AudioProtoPNet and BirdMAE show mixed behaviour with stronger overconfidence at low probabilities and underconfidence at higher probabilities. Second, the calibration is also influenced by the frequency of classes. We demonstrate a shift towards model overconfidence when aggregated across less-frequent classes. Surprisingly also, class-wise calibration appears better for less-frequent classes. This may point to potential bias when estimating calibration with limited samples. Investigating the reliability and bias of calibration metrics for rare classes remains an important direction for future work. Third, simple post hoc calibration can help, but its efficacy is model- and subset-dependent: global scaling tended to improve Perch v2 and ConvNeXtBS, while per-class Platt / temperature scaling often brought the largest gains, alongside occasional degradations that call for careful validation.

### v2

@@ 4.2 Temperature and Platt Scaling
Results for global TSare mixed. While ConvNeXtBS and Perch v2 show improvements in MCS on all test datasets ranging from 3.2% to 53%, AudioProtoPNet and BirdMAE exhibit more varied performance, with some datasets showing degradation in MCS by over 938%. On some datasets the improvements are substantial, with up to 98%. For AudioProtoPNet, TSseems to degrade calibration substantially on UHH and HSN and on PER (where the small relative change masks a sign flip from overconfident to underconfident). BirdMAE improves except for UHH. PSfollows the same trend as TS, with slightly better performance in most cases.
@@
The difference of MCS on the remaining test set (- first 10 min) compared to the full test set is marginal for all models. Comparing global and per-class scaling parameters is therefore valid. Per-class TSshows improvements for all models and datasets except for AudioProtoPNet on UHH, POW, and HSN, BirdMAE on POW and UHH and Perch v2 on POW and SNE. Here, also substantial degradations of up to -288% are observed. Improvements range up to 97%. Per-class PSimproves MCS for nearly all model–dataset combinations, with improvements up to 100%, and typically outperforms TS.
Both post hoc calibration methods demonstrate that substantial improvements to the deployment-specific calibration can be achieved with minor computational requirements. For example, AudioProtoPNet appears better calibrated than Perch v2 after per-class PS. However, no model or post hoc method consistently outperforms the others across all datasets, which reiterates that calibration and improvements using post hoc methods are domain-specific.
@@ 4.3 Implications for deployment
    When communicating uncertainty to end users (e.g., survey practitioners), reporting OCS/UCS alongside ECE or MCS helps reveal whether observed errors stem from systematic over- or underconfidence and whether that profile differs between common vs. rare species.
@@ 5 CONCLUSIONS
We presented a first comprehensive benchmark of uncertainty calibration for four multi-label deep learning bird sound classifiers across seven expert-annotated BirdSet test datasets and the POW validation set. Beyond reporting class-wise mean average precision (cmAP), we evaluated calibration holistically using reliability diagrams and threshold-free metrics: expected calibration error (ECE), miscalibration score (MCS), and the decomposition into over and underconfidence (OCS/UCS). We analysed calibration at three granularities—global, per-dataset, and per-class—and examined behaviour for common versus rare species. Finally, we assessed lightweight post hoc methods (temperature and Platt scaling) fitting global calibration parameters on the POW validation set and per-class parameters on the first 10 min of each test dataset.
Three findings stand out. First, calibration varies substantially across datasets and classes; aggregating across domains can mask systematic miscalibration. In our setting, Perch v2 and ConvNeXtBS are consistently underconfident (near-zero OCS, UCS-dominated MCS), whereas AudioProtoPNet and BirdMAE show mixed behaviour with stronger overconfidence at low probabilities and underconfidence at higher probabilities. Second, calibration is also influenced by the frequency of classes. We demonstrate a shift towards model overconfidence when aggregated across less-frequent classes. Also surprisingly, class-wise calibration appears better for less-frequent classes. This may point to potential bias when estimating calibration with limited samples. Investigating the reliability and bias of calibration metrics for rare classes remains an important direction for future work. Third, simple post hoc calibration can help, but its efficacy is model- and subset-dependent: global scaling tended to improve Perch v2 and ConvNeXtBS, while per-class Platt / temperature scaling often brought the largest gains, alongside occasional degradations that call for careful validation.